# STATIC : Surface Temporal Affine for TIme Consistency in Video Monocular Depth Estimation

### v1

@@ Abstract
{strip}[Uncaptioned image]
Figure 1: Results of the (a) Depth Anything v2 [40], (b) GEDepth [41] and proposed (c) STATIC from sequential frames. (c) improves ground continuity over the single-frame approach (a) and achieves better temporal consistency in object shape and depth compared to other video depth estimation method (b). For visual comparison, each method is rescaled.
@@ 1 Introduction
Depth estimation aims to generate a dense, pixel-level depth map from an RGB image, which is essential in applications such as autonomous driving, AR/VR, and robotics. Recently, transformer-based monocular depth estimation models  [27, 2, 43, 27] have demonstrated superior performance due to their robust generalization abilities, relying on large-scale datasets of paired single-view images and depth maps. In real-world applications like autonomous driving and robotics, depth maps are typically required from consecutive video frames. Monocular depth estimation can process video frames by predicting depth for each frame individually. However, it cannot consider inter-frame depth relationships. As illustrated in Figure 1 (a), per-frame predictions exhibit low inter-frame consistency. To handle these issues, several recent methods address temporal inconsistency by leveraging the global context inferred from multiple frames, enhancing connectivity across the temporal dimension [45, 23, 42]. By using multiple-frame inputs, the model improves inter-frame consistency by learning from diverse scenes. Others use explicit cues like optical flow [7, 38] or camera parameters [41, 3, 31] with high-quality motion estimation to enhance temporal consistency in depth predictions.
However, each previous temporal consistency method faces two major problems. Firstly, many multi-frame methods often focus on broad frame-to-frame changes rather than detailed local motions, making it harder for the model to capture subtle movements. Additionally, processing multiple frames simultaneously leads to inefficiencies and high memory consumption. Secondly, methods using explicit cues poorly perform with dynamic and irregular movements, leading to inaccurate motion information. Moreover, these methods struggle to independently capture the movements of both the foreground and background. This limitation often causes the edges of foreground elements to blur or lose clarity as the background shifts. As shown in Figure 1(b), unclear outlines can be observed for the same object between frames. Additionally, methods using explicit cues can become overly dependent on these additional cues, as mentioned in [17].
@@
As shown in Figure 2 (c), the proposed STATIC ensures spatial and temporal consistency across all areas. Our method was evaluated on two widely-used datasets: KITTI Eigen split [11], NYUv2 [30]. STATIC achieves state-of-the-art performance on both without additional information.
@@ 2.1 Monocular Depth Estimation
Monocular depth estimation, driven by deep learning, predicts depth from a single image as a practical alternative to multi-camera or radar setups. It mainly includes continuous regression for pixel-wise depth [44, 12, 46, 5, 29] and classification-based methods that use ordinal regression for depth ordering [4, 8, 19]. Recent integration of Transformer architectures [40, 27, 43] further enhances depth accuracy, with encoders capturing long-range spatial dependencies and decoders improving feature fusion for refined depth maps  [2]. Despite these advances, monocular depth estimation is fundamentally limited by its reliance on single-frame analysis. By relying on individual images, monocular methods miss temporal cues from consecutive frames, which are essential for capturing depth changes over time. This absence of temporal information makes it challenging for single-image monocular approaches to adapt to dynamic scene variations, such as moving objects or changes perspective. As a result, monocular depth models often produce less stable predictions in scenarios where scene depth fluctuates. This instability limits their effectiveness in video applications where consistent depth tracking is essential.
@@ 2.2 Video Depth Estimation
Video depth estimation builds on monocular depth prediction by incorporating temporal consistency and motion cues across frames, enhancing stability and accuracy over time. Unlike single-frame approaches, it captures temporal dependencies to resolve frame-to-frame inconsistencies in dynamic scenes. Recent advancements leverage RNNs [7, 23, 45, 42] and optical flow [10, 32, 13] for improved temporal alignment. RNNs use memory to link depth predictions across frames and enhance stability, while optical flow provides explicit motion cues, aligning depth with object dynamics. Transformers have also been adopted to capture both spatial and temporal dependencies, improving consistency over time. More recently, diffusion models [24, 39], supported by larger datasets, have been explored. however, they mainly enhance data diversity rather than directly tackling the unique challenges of video depth consistency. Despite these advances, challenges remain, particularly the high memory demands, dependencies, and computational costs required to process multiple frames for temporal consistency. These limitations highlight the need for further research to achieve more stable and efficient video depth estimation.
2.3 Surface Normal
Surface normals, describing the orientation of surfaces within a scene, are closely related to depth and provide complementary geometric information crucial for maintaining spatial and temporal consistency. This relationship is strong enough that simple operators, like Sobel filters [14], can effectively capture surface normal cues from depth variations, offering a fundamental basis for depth estimation models to understand structural changes across frames. By analyzing directional variations, surface normals assist in identifying changes in scene geometry, allowing depth models to adapt more robustly to transformations without relying on external motion data. This approach enhances stability in dynamic environments, capturing depth details more reliably by leveraging the inherent geometry of the scene.
@@ 3.1 Overall Architecture
Figure 2 shows the overall architecture of STATIC. Our model uses only two frames, $I_{t}\in\mathbb{R}^{3\times H\times W}$, where $t\in[0,1]$, as input. $I_{t}$ is passed through an encoder to obtain the encoder feature. These features are shared by both the surface normal decoder and the depth embedder. The depth embedder processes these features to produce the depth context feature $Z^{d}_{t}\in\mathbb{R}^{C^{d}\times\frac{H}{8}\times\frac{W}{8}}$. Additionally, the encoder feature is passed through a simple surface normal decoder. This decoder generates both the surface normal image $N_{t}\in\mathbb{R}^{3\times\frac{H}{8}\times\frac{W}{8}}$ and the surface normal features $Z^{n}_{t}\in\mathbb{R}^{C^{n}\times\frac{H}{8}\times\frac{W}{8}}$, which is a multi-scale normal decoder feature. $N_{t}$ is utilized to generate a difference mask, which serves to separate specific areas. STATIC obtains temporal consistency by employing both the Masked Static (MS) module and the Surface Normal Similarity (SNS) module. The MS module takes depth features and a difference mask as inputs. It applies masking to each frame’s depth feature to generate temporal consistency information for the static area. Additionally, the SNS module takes depth features, surface normal features, and a difference mask as inputs. It ensures alignment and maintains temporal consistency in the dynamic area. Finally, the independently learned SNS and MS module features are refined and connected to obtain a video feature with temporal consistency. This video feature is concatenated with the output from the depth embedder. The depth head then performs upsampling to generate consecutive depth maps for both frames.
@@ 3.2 Difference Mask
$M^{d}=\begin{cases}1\qquad&\text{if }\quad\sigma^{2}_{total}>(\mu_{\sigma^{2}}+\alpha)\\ 0\qquad&\text{otherwise}\end{cases}$ (2)
@@ 3.5 Loss function
Following previous works [44, 15], we use a scaled version of the Scale-Invariant loss (SILog) [6] to train the depth map. In addition, we use Mean Squared Error (MSE) to supervise the surface normal. Similarly, MSE is also employed in the process of generating the refined difference mask $M^{l}$. The $M^{l}$ generation loss is defined as follows:
@@
where $L_{Depth}$ is SILog, while $L_{Normal}$ and $L_{Mask}$ are MSE. Following [15], we use the SILog parameter $\lambda=0.85$. In addition, we set $\alpha=10$.
@@ 4 Experiments
Method Frame C O M Abs Rel $\downarrow$ Sq Rel $\downarrow$ RMSE $\downarrow$ $\delta<1.25$uparrow $\delta<1.25^{2}$uparrow $\delta<1.25^{3}$uparrow
PEM  [16] SF - - - 0.068 0.221 2.127 0.958 0.993 0.9983
AdaBins  [4] SF - - - 0.058 0.190 2.360 0.964 0.995 0.9993
BinsFormer  [19] SF - - - 0.058 0.190 2.336 0.964 0.996 0.9994
DepthFormer  [1] SF - - - 0.053 0.187 2.285 0.970 0.996 0.9994
PixelFormer  [2] SF - - - 0.052 0.152 2.093 0.975 0.997 0.9994
GEDepth  [41] SF $\checkmark$ - - 0.049 0.143 2.048 0.976 0.997 0.9994
NeuralRGB  [20] MF $\checkmark$ - - 0.100 - 2.829 - - -
ST-CLSTM  [45] MF - - $\checkmark$ 0.101 - 4.137 0.890 0.970 0.9890
FlowGRU  [7] MF - $\checkmark$ $\checkmark$ 0.112 0.070 4.260 0.936 0.983 0.9930
Flow2Depth  [38] MF $\checkmark$ $\checkmark$ - 0.109 - 4.284 0.910 0.980 0.9900
RDE-MV  [23] MF - - $\checkmark$ 0.111 0.821 4.650 0.898 0.972 0.9890
FMNet  [34] MF - - - 0.069 0.342 3.340 0.946 0.986 0.9960
ManyDepth-FS  [36] MF $\checkmark$ - - 0.053 0.243 2.248 0.975 0.997 0.9994
TC-Depth-FS  [28] MF - - - 0.059 0.249 3.280 0.947 0.985 0.9940
MAMo  [42] MF - $\checkmark$ $\checkmark$ 0.049 0.130 1.989 0.977 0.998 0.9995
@@
4.1 Implementation Details
The proposed method is implemented using the AdamW [22] optimizer with $\beta$ parameters of 0.9 and 0.999, a batch size of 2 with 2 frames per batch, and a weight decay set to $10^{-2}$. We train the model over 20 epochs on both the KITTI and NYUv2 datasets, starting from an initial learning rate of $4\times 10^{-5}$, which linearly decays to $4\times 10^{-6}$ over the training process. With four NVIDIA 4090 GPUs, each epoch takes approximately 60 minutes. The encoder backbone is initialized with pre-trained Swinv2-L [21] weights. During testing, final depth predictions are obtained by averaging the outputs from both the original and mirrored inputs. We employ a training and testing protocol similar to those used in [2, 15].
@@ 4.2 Evaluation Metrics
Additionally, we use a metric $s$ from Li et al. [18] for evaluating temporal consistency. This metric is as follows:
@@
where $D_{t}$ is the predicted depth, $D_{t}^{w}$ is the warped depth from $D_{t-1}$, and $K^{t}$ is a depth validity mask. We use Flowformer [13] as the optical flow model for warping. We use this metric to present the temporal consistency comparison results in Table 3.
@@
4.3 Datasets
KITTI Eigen: The KITTI dataset [11] is among the most frequently utilized benchmarks for outdoor depth estimation. In our approach, we adopt the Eigen split for training and testing, which includes 23,488 training images and 697 test images. Video sequences corresponding to these training and test images are utilized, with each video frame having a resolution of 375×1241 pixels. For evaluating the test set, we apply the crop defined by Garg et al. [9], and depth estimation is performed up to a distance of 80 meters.
NYU Depth v2: NYU v2 [30] is a well-known indoor dataset, providing 120,000 RGB and depth image pairs (480×640 resolution) collected as video sequences across 464 indoor environments. For video-based evaluation, we adapted the dataset by employing the test approach adopted by [35]. Specifically, we used 249 scenes from the original 464 scenes, comprising pairs of RGB and sync depth images  [15] for training. The remaining 215 scenes containing 654 images were used for testing. Each depth map is limited to a maximum range of 10 meters, and we apply center cropping as suggested by Eigen et al. [6].
@@ 4.4 Result
Results on KITTI: Table 1 presents our results on the KITTI dataset, where STATIC demonstrates robust performance with minimal input requirements. We compare our approach against various state-of-the-art multi-frame and single-frame methods. Even without camera parameters, additional memory, or optical flow, our model achieves competitive results across all metrics. Our model’s depth embedder is based on PixelFormer [2] and outperforms the baseline, highlighting the effectiveness of the temporal consistency module. Furthermore, our approach improves RMSE by approximately 0.60% over the recently proposed MAMo [42], even though MAMo uses more frames. These results indicate that our model effectively preserves consistency in independent areas, even in the presence of substantial movement within outdoor environments.
Type Method $\delta<1.25$uparrow Abs Rel $\downarrow$
SF Midas-v2.1-Large [26] 0.910 0.095
DPT-Large  [27] 0.928 0.084
WSVD  [33] 0.768 0.164
ST-CLSTM  [45] 0.833 0.131
MF DeepV2D  [31] 0.924 0.082
FMNet  [34] 0.832 0.134
VITA  [37] 0.922 0.092
MAMo  [42] 0.919 0.094
@@
Results on NYU v2: Table 2 presents our results on the NYU v2 dataset. Following the evaluation procedure of [2] and without additional training data, our method achieves state-of-the-art performance on the $\delta<1.25$ metric. Specifically, our model demonstrates a $1.63\%$ improvement over the recently proposed MAMo method on $\delta<1.25$ and a $0.65\%$ gain over the previous state-of-the-art. Furthermore, our model shows effective improvements on the Abs Rel metric compared to other multi-frame methods.

### v2

@@ 1 Introduction
Depth estimation aims to generate a dense, pixel-level depth map from an RGB image, which is essential in applications such as autonomous driving, AR/VR, and robotics. Recently, transformer-based monocular depth estimation models  [ranftl2021vision, agarwal2023attention, yuan2022neural, ranftl2021vision] have demonstrated superior performance due to their robust generalization abilities, relying on large-scale datasets of paired single-view images and depth maps. In real-world applications like autonomous driving and robotics, depth maps are typically required from consecutive video frames. Monocular depth estimation can process video frames by predicting depth for each frame individually. However, it cannot consider inter-frame depth relationships. As illustrated in Figure LABEL:fig:fig1 (a), per-frame predictions exhibit low inter-frame consistency. To handle these issues, several recent methods address temporal inconsistency by leveraging the global context inferred from multiple frames, enhancing connectivity across the temporal dimension [zhang2019exploiting, patil2020don, yasarla2023mamo]. By using multiple-frame inputs, the model improves inter-frame consistency by learning from diverse scenes. Others use explicit cues like optical flow [eom2019temporally, xie2020video] or camera parameters [yang2023gedepth, azinovic2022neural, teed2018deepv2d] with high-quality motion estimation to enhance temporal consistency in depth predictions.
However, each previous temporal consistency method faces two major problems. Firstly, many multi-frame methods often focus on broad frame-to-frame changes rather than detailed local motions, making it harder for the model to capture subtle movements. Additionally, processing multiple frames simultaneously leads to inefficiencies and high memory consumption. Secondly, methods using explicit cues poorly perform with dynamic and irregular movements, leading to inaccurate motion information. Moreover, these methods struggle to independently capture the movements of both the foreground and background. This limitation often causes the edges of foreground elements to blur or lose clarity as the background shifts. As shown in Figure LABEL:fig:fig1(b), unclear outlines can be observed for the same object between frames. Additionally, methods using explicit cues can become overly dependent on these additional cues, as mentioned in [lee2024video].
@@
As shown in Figure LABEL:fig:fig1 (c), the proposed STATIC ensures spatial and temporal consistency across all areas. Our method was evaluated on two widely-used datasets: KITTI Eigen split [geiger2012we], NYUv2 [silberman2012indoor]. STATIC achieves state-of-the-art performance on both without additional information.
@@ 2.1 Monocular Depth Estimation
Monocular depth estimation, driven by deep learning, predicts depth from a single image as a practical alternative to multi-camera or radar setups. It includes continuous regression for pixel-wise depth [yuan2022new, godard2017unsupervised, zhu1232020mda, cai2021x, shi2023ega] and classification-based methods using ordinal regression for depth ordering [bhat2021adabins, fu2018deep, li2024binsformer]. Recent integration of Transformer architectures [yang2024depth, ranftl2021vision, yuan2022neural] has further improved accuracy, with encoders capturing long-range dependencies and decoders enhancing feature fusion for refined maps [agarwal2023attention]. Despite these advances, monocular depth estimation remains limited by single-frame analysis, missing temporal cues essential for capturing depth changes. Without temporal information, it struggles to adapt to dynamic scenes with moving objects or perspective shifts, often producing unstable predictions when scene depth fluctuates. This instability reduces effectiveness in video applications where consistent depth tracking is crucial.
@@ 2.2 Video Depth Estimation
Video depth estimation builds on monocular depth prediction by incorporating temporal consistency and motion cues across frames, enhancing stability and accuracy over time. Unlike single-frame approaches, it captures temporal dependencies to resolve frame-to-frame inconsistencies in dynamic scenes. Recent advancements leverage RNNs [eom2019temporally, patil2020don, zhang2019exploiting, yasarla2023mamo] and optical flow [garrepalli2023dift, teed2020raft, huang2022flowformer] for improved temporal alignment. RNNs use memory to link depth predictions across frames and enhance stability, while optical flow provides explicit motion cues, aligning depth with object dynamics. Transformers have also been adopted to capture both spatial and temporal dependencies, improving consistency over time. More recently, diffusion models [patni2024ecodepth, fu2024geowizard, ke2024repurposing], supported by larger datasets, have been explored. However, they mainly enhance data diversity rather than directly tackling the unique challenges of video depth consistency. Despite these advances, challenges remain, particularly the high memory demands and computational costs required to process depth estimation. These limitations highlight the need for further research to achieve more stable and efficient video depth estimation.
2.3 Surface Normal-based Approaches
Previous methods [long2020occlusion, long2021adaptive] have exploited surface normals to enhance the geometric quality of depth by preserving structures such as corners, boundaries, and planes. Surface normals from local depth gradients remain stable across frames, as gradient directions do not change significantly when large structures are preserved. Even when depth values show low correlation across frames, surface normals remain consistent. However, prior works overlooked this property and focused mainly on sharp edges. Our contribution is to explicitly leverage inter-frame normal consistency in a video-based method, for the first time. Using this property, we compute inter-frame differences by analyzing the variation and dispersion of surface normal directions at each pixel, thereby identifying regions that differ in the subsequent frame. In these regions, our proposed SNS and MS modules enforce consistency, enabling temporally consistent video depth prediction.
@@ 3 Method
3.1 Preliminaries
Surface normals, which describe the orientation of surfaces within a scene, are closely related to depth and provide complementary geometric information crucial to maintaining spatial and temporal consistency. This relationship is strong enough that simple operators, like Sobel filters [kittler1983accuracy], can effectively capture surface normal cues from depth variations, offering a fundamental basis for depth estimation models to understand structural changes across frames. By analyzing directional variations, surface normals assist in identifying changes in scene geometry, allowing depth models to adapt more robustly to transformations without relying on external motion data. This approach enhances stability in dynamic environments, capturing depth details more reliably by leveraging the inherent geometry of the scene.
@@ 3.2 Overall Architecture
Figure 2 shows the overall architecture of STATIC. Our model uses only two frames, $I_{t}\in\mathbb{R}^{3\times H\times W}$, where $t\in[0,1]$, as input. $I_{t}$ is passed through an encoder to obtain the encoder feature. These features are shared by both the surface normal decoder and the depth embedder. The depth embedder processes these features to produce the depth context feature $Z^{d}_{t}\in\mathbb{R}^{C^{d}\times\frac{H}{8}\times\frac{W}{8}}$. Additionally, the encoder feature is passed through a simple surface normal decoder. This decoder generates both the surface normal image $N_{t}\in\mathbb{R}^{3\times\frac{H}{8}\times\frac{W}{8}}$ and the surface normal features $Z^{n}_{t}\in\mathbb{R}^{C^{n}\times\frac{H}{8}\times\frac{W}{8}}$, which is a multi-scale normal decoder feature. $N_{t}$ is utilized to generate a difference mask, which serves to separate regions of structural change. STATIC obtains temporal consistency by employing both the Masked Static (MS) module and the Surface Normal Similarity (SNS) module. The MS module takes depth features and a difference mask as inputs. It applies masking to each frame’s depth feature to generate temporal consistency information for the static area. Additionally, the SNS module takes depth features, surface normal features, and a difference mask as inputs. It ensures alignment and maintains temporal consistency in the dynamic area. Finally, the independently learned SNS and MS module features are refined and connected to obtain a video feature with temporal consistency. This video feature is concatenated with the output from the depth embedder. The depth head then performs upsampling to generate consecutive depth maps for both frames.
@@ 3.3 Difference Mask
[…]
@@ 3.6 Loss function
Following previous works [yuan2022new, lee2019big], we use a scaled version of the Scale-Invariant loss (SILog) [eigen2014depth] to train the depth map. In addition, we use Mean Squared Error (MSE) to supervise the surface normal. Similarly, MSE is also employed in the process of generating the refined difference mask $M^{l}$. The $M^{l}$ generation loss is defined as follows:
@@
where $L_{Depth}$ is SILog, while $L_{Normal}$ and $L_{Mask}$ are MSE. Following [lee2019big], we use the SILog parameter $\lambda=0.85$. In addition, we set $\alpha=10$.
@@ 4 Experiments
Method Frame C O M Abs Rel $\downarrow$ Sq Rel $\downarrow$ RMSE $\downarrow$ $\delta<1.25$ $\uparrow$ $\delta<1.25^{2}$ $\uparrow$ $\delta<1.25^{3}$ $\uparrow$
PEM  [lee2022edgeconv] SF - - - 0.068 0.221 2.127 0.958 0.993 0.9983
AdaBins  [bhat2021adabins] SF - - - 0.058 0.190 2.360 0.964 0.995 0.9993
BinsFormer  [li2024binsformer] SF - - - 0.058 0.190 2.336 0.964 0.996 0.9994
DepthFormer  [agarwal2022depthformer] SF - - - 0.053 0.187 2.285 0.970 0.996 0.9994
PixelFormer  [agarwal2023attention] SF - - - 0.052 0.152 2.093 0.975 0.997 0.9994
GEDepth  [yang2023gedepth] SF $\checkmark$ - - 0.049 0.143 2.048 0.976 0.997 0.9994
NeuralRGB  [liu2019neural] MF $\checkmark$ - - 0.100 - 2.829 - - -
ST-CLSTM  [zhang2019exploiting] MF - - $\checkmark$ 0.101 - 4.137 0.890 0.970 0.9890
FlowGRU  [eom2019temporally] MF - $\checkmark$ $\checkmark$ 0.112 0.070 4.260 0.936 0.983 0.9930
Flow2Depth  [xie2020video] MF $\checkmark$ $\checkmark$ - 0.109 - 4.284 0.910 0.980 0.9900
RDE-MV  [patil2020don] MF - - $\checkmark$ 0.111 0.821 4.650 0.898 0.972 0.9890
FMNet  [wang2022less] MF - - - 0.069 0.342 3.340 0.946 0.986 0.9960
ManyDepth-FS  [watson2021temporal] MF $\checkmark$ - - 0.053 0.243 2.248 0.975 0.997 0.9994
TC-Depth-FS  [ruhkamp2021attention] MF - - - 0.059 0.249 3.280 0.947 0.985 0.9940
MAMo  [yasarla2023mamo] MF - $\checkmark$ $\checkmark$ 0.049 0.130 1.989 0.977 0.998 0.9995
@@ 4.1 Evaluation Metrics
Additionally, we use a metric $s$ from Li et al. [li2021enforcing] for evaluating temporal consistency. This metric is as follows:
@@
where $D_{t}$ is the predicted depth, $D_{t}^{w}$ is the warped depth from $D_{t-1}$, and $K^{t}$ is a depth validity mask. We use Flowformer [huang2022flowformer] as the optical flow model for warping. We use this metric to present the temporal consistency comparison results in Table 3.
4.2 Datasets
KITTI Eigen: The KITTI dataset [geiger2012we] is among the most frequently utilized benchmarks for outdoor depth estimation. In our approach, we adopt the Eigen split for training and testing, which includes 23,488 training images and 697 test images. Video sequences corresponding to these training and test images are utilized, with each video frame having a resolution of 375×1241 pixels. For evaluating the test set, we apply the crop defined by Garg et al. [garg2016unsupervised], and depth estimation is performed up to a distance of 80 meters.
NYU Depth v2: NYU v2 [silberman2012indoor] is a well-known indoor dataset, providing 120,000 RGB and depth image pairs (480×640 resolution) collected as video sequences across 464 indoor environments. For video-based evaluation, we adapted the dataset by employing the test approach adopted by [wang2023neural]. Specifically, we used 249 scenes from the original 464 scenes, comprising pairs of RGB and sync depth images  [lee2019big] for training. The remaining 215 scenes containing 654 images were used for testing. Each depth map is limited to a maximum range of 10 meters, and we apply center cropping as suggested by Eigen et al. [eigen2014depth].
@@ 4.3 Result
Results on KITTI: Table 1 presents our results on the KITTI dataset, where STATIC demonstrates robust performance with minimal input requirements. We compare our approach against various state-of-the-art multi-frame and single-frame methods. Even without camera parameters, additional memory, or optical flow, our model achieves competitive results across all metrics. Our model’s depth embedder is based on PixelFormer [agarwal2023attention] and outperforms the baseline, highlighting the effectiveness of the temporal consistency module. Furthermore, our approach improves RMSE by approximately 0.60% over the recently proposed MAMo [yasarla2023mamo], even though MAMo uses more frames. These results indicate that our model effectively preserves consistency in independent areas, even in the presence of substantial movement within outdoor environments.
Type Method $\delta<1.25$ $\uparrow$ Abs Rel $\downarrow$
SF Midas-v2.1-Large [ranftl2020towards] 0.910 0.095
DPT-Large  [ranftl2021vision] 0.928 0.084
WSVD  [wang2019web] 0.768 0.164
ST-CLSTM  [zhang2019exploiting] 0.833 0.131
MF DeepV2D  [teed2018deepv2d] 0.924 0.082
FMNet  [wang2022less] 0.832 0.134
VITA  [xian2023vita] 0.922 0.092
MAMo  [yasarla2023mamo] 0.919 0.094
@@
Results on NYU v2: Table 2 presents our results on the NYU v2 dataset. Following the evaluation procedure of [agarwal2023attention] and without additional training data, our method achieves state-of-the-art performance on the $\delta<1.25$ metric. Specifically, our model demonstrates a $1.63\%$ improvement over the recently proposed MAMo method on $\delta<1.25$ and a $0.65\%$ gain over the previous state-of-the-art. Furthermore, our model shows effective improvements on the Abs Rel metric compared to other multi-frame methods.